# Multigraph Message Passing with Bi-Directional Multi-Edge Aggregations

### v1

@@ 6 Discussion and Conclusion
Our experiments on financial transaction datasets show that MEGA-GNN significantly outperforms or matches the accuracy of state-of-the-art multigraph GNN models. Notably, when detecting illicit transactions on the synthetic AML datasets of Altman et al. (2023), we achieve improvements of up to $13.31\%$ in minority-class F1 scores over the method of Egressy et al. (2024), demonstrating the effectiveness of MEGA-GNN. When detecting phishing accounts on the Ethereum transaction dataset of (Chen et al. (2021)), we achieve improvements of around $20\%$ in minority-class F1 scores over the method of Sotiropoulos et al. (2023), while slightly improving the results of Egressy et al. (2024).
@@ 7 Reproducibility Statement
To ensure the reproducibility of our research, we will open-source our codes in a publicly accessible repository. The experiments conducted as part of this work used only publicly available datasets.
@@ 8 Ethics Statement
Appendix A Appendix
@@
Since we are assuming the GNN is permutation-equivariant, the latent node features should remain the same after permuting the port numbers:
@@
This implies that for all possible permutations $\sigma$, the latent features of the nodes should be invariant.
Figure 6: Counterexample for the permutation-equivariance of Multi-GNN Egressy et al. (2024). The left panel shows the graph ${\mathcal{G}}$ with one permutation of the port numbering, while the right panel illustrates a different permutation of the assigned port numbers. Notice how the node embeddings change when computed using Algorithm 1 from Egressy et al. (2024), highlighting that different port number assignments lead to different embeddings, thus violating permutation-equivariance
Let the GNN, $f$, mimics the Algorithm 1 from Egressy et al. (2024), which computes unique node IDs based on port numbers. If we apply this GNN, $f$, to a Star Graph with order $n>3$ with different port number permutations, resulting latent node features must be distinct. Formally, applying $f$ to a the graph with port numbering $p$ gives the latent node features $\mathbf{X^{(l)}}$, and applying $f$ to the graph with a permuted port numbering $p_{\sigma}$, gives a different set of node embeddings $\mathbf{\hat{X}}^{(l)}$:
@@
This result directly contradicts our earlier assumption that the model is permutation-equivariant with respect to node ordering, preserves the identity of node outputs. Specifically, if the GNN were permutation-equivariant, we would expect $\mathbf{X}^{(l)}=\mathbf{\hat{X}}^{(l)}$, but instead the permutation of port numbers leads to different output. Illustration of this contradictory example is given in Figure 6.
Therefore, we conclude that the arbitrary assignment of port numbers is not permutation-equivariant with respect to node orderings.
@@
Proof: The proposed message passing layer performs two aggregations over the neighborhood of a target node $j$. The first is the multi-edge aggregation, in which the latent features of the parallel edges are aggregated at artificial nodes. The one-hop neighborhood of artificial node $v_{ij}^{art}$ is given by the multiset $ME_{ij}$, the cardinality of which is $P_{ij}$:
$\begin{split}\mathbf{E}_{ME_{ij}}=\{\mathbf{e}_{ijp}\ |\ p\in\{1..P_{ij}\}\}.\end{split}$ (15)
Let’s define a permutation-invariant multi-edge aggregation function $\psi$, which operates on the neighborhood of an artificial node (the parallel edges):
@@
Since $\psi$ is a permutation-invariant function, for any permutation function $\rho$ acting on the neighboring edges (parallel edges) of the artificial node, we have $\psi(\rho\cdot\mathbf{E}_{ME_{ij}})=\rho\cdot\psi(\mathbf{E}_{ME_{ij}})$.
@@
Similarly, let’s define a permutation-invariant node-level aggregation function $\phi$, which operates over the neighborhood of the target nodes:
@@
Since $\phi$ is a permutation-invariant function, for any permutation function $\pi$ acting on the neighboring edges of a target node $i$, we have $\phi(\pi\cdot(\mathbf{x}_{j},\mathbf{X}_{N_{in}(j)}))=\pi\cdot\phi((\mathbf{x}_{j},\mathbf{X}_{N_{in}(j)}))$.
In our framework, $\psi$ performs multi-edge aggregation, handling parallel edges, while $\phi$ is responsible for node-level aggregation, combining messages from distinct neighbors of the target node. The two-stage aggregation scheme integrates $\psi$ and $\phi$ within a single message passing layer. Since the composition of permutation-invariant functions remains permutation-invariant, our message passing layer ($\psi\circ\phi$) is invariant to the permutations of neighboring nodes and edges ($\rho\circ\pi$). Unlike simple graphs, node permutations do not directly imply edge permutations due to the presence of parallel edges. Thus, we explicitly define the permutation of parallel edges, $\rho$, ensuring that our message passing layer remains permutation-invariant to both nodes and edges in the neighborhood of the target node.
Finally, as demonstrated by Bronstein et al. (2021), the composition of permutation-invariant layers ($f=\psi\circ\phi\circ\psi\circ\phi\cdots$) allows the construction of functions $f$ that are equivariant to symmetry group actions. In the multigraph domain, this symmetry group includes permutations of both nodes and edges, as node permutations do not directly induce edge permutations due to the presence of parallel edges. The overall permutation equivariance of the MEGA-GNN model follows from the fact that each permutation-invariant message passing layer operates independently on each node’s neighborhood, regardless of the ordering of nodes or edges. Specifically, for any permutation $g\in\sum_{n}$ acting on the set of node and edge indices, the model’s output satisfies $f(g\cdot X)=g\cdot f(X)$.
@@
Egressy et al. (2024) showed that a GNN can mimic a Breadth-First Search (BFS) algorithm to compute unique node IDs given pre-computed port IDs for the edges. We follow the same BFS-based approach and derive unique node ids without relying on pre-computed port IDs. Instead of the pre-computed port ids, we use the unique edge labels provided by $L(e)$ to guide the node ID assignment process. As in Egressy et al. (2024), we use the Universal Approximation Theorem Hornik et al. (1989) for MLPs, to avoid explicit construction of the MEGA-GNN layers. We also assume that the MEGA-GNN aggregates the multi-edges by computing their minimum, which is followed by a node-level aggregation, where an MLP is applied element-wise to the incoming messages, followed by another minimum computation.
Following the approach of Egressy et al. (2024), the node ID assignment algorithm starts from a root node (also called the ego node) and assigns IDs to all the other nodes connected to it via message passing . We are not going to reiterate the setup of the entire proof and focus on the differences. Namely, instead of pre-computing port-IDs, we use an edge labeling function, which assumes a strict total ordering among the edges based on the original edge features. In addition, the multi-edges are aggregated by selecting the edge with the minimum label.
Our MEGA-GNN model, which mimics Algorithm 1, assigns ids to each node connected to the root node. What remains to be shown is that those assigned node IDs are unique. First, note that nodes at different distances from the root cannot end up with the same node ID. A node at distance $k$ will receive its first proposal in round $k$ and, therefore, it will have an ID with exactly $k+1$ digits. Furthermore, an inductive argument shows that active nodes (nodes at the same distance) cannot have the same node IDs. Certainly, this is true at the start when $X=\{r\}$. Now assuming all active nodes from the previous round $(k-1)$ had distinct node IDs, then the only way two active nodes (in round $k$) can have the same ID is if they accept a proposal from the same neighboring node. This is because, based on the induction hypothesis, proposals from different nodes will already differ in their first $k-1$ digits. If two active nodes accepted a proposal from the same node, then they would have received different edge labels —a strict total ordering among the edges enables assignment of distinct edge labels. In addition, because $m$ is added to all incoming labels, incoming labels cannot be the same as the outgoing labels. Therefore the active nodes always accept unique proposals.
@@ B.1 Hyperparameters
Table 5: Hyperparameter settings for AML and ETH
@@ B.2 Multi-Edge Aggregation
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{MLP}\bigg{(}\text{SUM}\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\bigg{)}.\end{split}$ (19)
@@
In the PNA aggregation we combined [’mean’, ’min’, ’max’, ’std’] statistics in combination with scalars [’identity’, ’amplification’, ’attenuation’] around the neighborhood of each artificial node, $N_{in}(v_{ij}^{art})=ME_{ij}$.
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{MLP}\bigg{(}\text{PNA}\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\bigg{)}.\end{split}$ (20)
@@
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{GenAgg}\bigg{(}\text{SUM}\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\bigg{)}.\end{split}$ (21)
@@ Appendix C Additional Performance Metrics for AML Results
Table 6: AML edge classification task results for Small datasets.
Small_LI Small_HI
@@
Table 7: AML edge classification task results for Medium datasets.
Medium_LI Medium_HI
@@ Appendix E Ablation Experiment
In this section, we perform additional experiments on the baseline method, Multi-GNN, to investigate the effect of randomly assigning port-ids. The results, presented in Table 8, were obtained using the same set of hyperparameters as those used for MEGA-GIN and MEGA-PNA, applied to Multi-GIN and Multi-PNA, respectively.
Ablation Small_HI Small_LI ETH-Kaggle Medium_HI Medium_LI
Multi-GIN 62.65 $\pm$ 1.73 32.21 $\pm$ 0.99 51.34 $\pm$ 3.92 67.72 $\pm$ 0.93 31.24 $\pm$ 2.12
Multi-GIN (permuted) 61.36 $\pm$ 4.17 29.00 $\pm$ 3.49 49.13 $\pm$ 3.72 67.58 $\pm$ 1.96 30.19 $\pm$ 1.49
Multi-PNA 67.35 $\pm$ 2.89 35.39 $\pm$ 3.93 64.61 $\pm$ 1.40 71.93 $\pm$ 1.54 43.81 $\pm$ 0.51
Multi-PNA (permuted) 66.58 $\pm$ 1.86 33.05 $\pm$ 3.84 65.01 $\pm$ 0.97 73.41 $\pm$ 1.38 42.61 $\pm$ 1.21
Table 8: Multi-GNN ablations: The port-id assignment in Multi-GNN follows a strict total ordering among edges. In the permuted ablations, port-id assignment is randomized to evaluate the impact of edge ordering on model performance.
Randomly permuting port-id assignments leads to a slight performance drop for both models, with Multi-GIN experiencing a more significant decline compared to Multi-PNA. Notably, Multi-PNA demonstrates greater robustness to permuted port-ids. Interestingly, on the ETH-Kaggle dataset, permuting the port-ids results in a small performance improvement for the PNA-based model, suggesting that the permutation equivariance property is less critical for this particular dataset.

### v2

@@ 6 Discussion and Conclusion
Our experiments on financial transaction datasets show that MEGA-GNN significantly outperforms or matches the accuracy of state-of-the-art multigraph GNN models. Notably, when detecting illicit transactions on the synthetic AML datasets of Altman et al. (2023), we achieve improvements of up to $13.31\%$ in minority-class F1 scores over the state-of-the-art method (Egressy et al. (2024)), demonstrating the effectiveness of MEGA-GNN. When detecting phishing accounts on the Ethereum transaction dataset (Chen et al. (2021)), we achieve improvements of around $20\%$ in minority-class F1 scores over the method of Sotiropoulos et al. (2023), while slightly improving the results of Egressy et al. (2024).
@@ 7 Reproducibility Statement
To ensure the reproducibility of our research, we open-sourced our codes in a publicly accessible repository. The experiments conducted as part of this work used only publicly available datasets.
@@ 8 Ethics Statement
Acknowledgments
The support of the Swiss National Science Foundation (project number: 212158) for this work is gratefully acknowledged.
Appendix A Proofs
Figure 6: Counterexample for the permutation equivariance of Multi-GNN Egressy et al. (2024). The left panel shows the graph ${\mathcal{G}}$ with one permutation of the port numbering, while the right panel illustrates a different permutation of the assigned port numbers. Notice how the node embeddings change when computed using Algorithm 1 from Egressy et al. (2024), highlighting that different port number assignments lead to different embeddings, thus violating permutation equivariance
@@ Appendix A Proofs
Since we are assuming the GNN is permutation-equivariant, the latent node features of node $i$ should remain the same after permuting the port numbers:
@@
Let the GNN, $f$, mimics the Algorithm 1 from Egressy et al. (2024), which computes unique node IDs based on port numbers. If we apply this GNN, $f$, to a Star Graph with order $n>3$ with different permutations of port numbering, resulting latent node features must be distinct. Formally, applying $f$ to a the graph with port numbering $p$ gives the latent node features $\mathbf{X^{(l)}}$, and applying $f$ to the graph with a permuted port numbering $p_{\sigma}$, gives a different set of node embeddings $\mathbf{\hat{X}}^{(l)}$:
@@
This result directly contradicts our earlier assumption that the model is permutation equivariant with respect to node orderings. Specifically, if the GNN were permutation equivariant, we would expect $\mathbf{X}^{(l)}=\mathbf{\hat{X}}^{(l)}$, but instead the permutation of port numbers leads to different output. Illustration of this contradictory example is given in Figure 6.
Therefore, we conclude that the arbitrary assignment of port numbers is not permutation equivariant with respect to node orderings.
@@
Proof: The proposed message passing layer performs two aggregations over the neighborhood of a target node $j$. The first is the multi-edge aggregation, in which the latent features of the parallel edges are aggregated at artificial nodes. The one-hop neighborhood of artificial node $v_{ij}^{art}$ is given by the multiset $ME_{ij}$, with cardinality $P_{ij}$:
$\begin{split}\mathbf{E}_{ME_{ij}}=\{\!\{\ \mathbf{e}_{ijp}\ |\ p\in\{1..P_{ij}\}\ \}\!\}.\end{split}$ (15)
Let’s define a permutation invariant multi-edge aggregation function $\psi$, which operates on the neighborhood of an artificial node (the parallel edges):
@@
Since $\psi$ is a permutation invariant function, for any permutation function $\rho$ acting on the neighboring edges (parallel edges) of the artificial node, we have $\psi(\rho\cdot\mathbf{E}_{ME_{ij}})=\rho\cdot\psi(\mathbf{E}_{ME_{ij}})$.
@@
Similarly, let’s define a permutation invariant node-level aggregation function $\phi$, which operates over the neighborhood of the target nodes:
@@
Since $\phi$ is a permutation invariant function, for any permutation function $\pi$ acting on the neighboring edges of a target node $i$, we have $\phi(\pi\cdot(\mathbf{x}_{j},\mathbf{X}_{N_{in}(j)}))=\pi\cdot\phi((\mathbf{x}_{j},\mathbf{X}_{N_{in}(j)}))$.
In our framework, $\psi$ performs multi-edge aggregation, handling parallel edges, while $\phi$ is responsible for node-level aggregation, combining messages from distinct neighbors of the target node. The two-stage aggregation scheme integrates $\psi$ and $\phi$ within a single message passing layer. Since the composition of permutation invariant functions remains permutation invariant, our message passing layer ($\psi\circ\phi$) is invariant to the permutations of neighboring nodes and edges ($\rho\circ\pi$). Unlike simple graphs, node permutations do not directly imply edge permutations due to the presence of parallel edges. Thus, we explicitly define the permutation of parallel edges, $\rho$, ensuring that our message passing layer remains permutation-invariant to both nodes and edges in the neighborhood of the target node.
Finally, as demonstrated by Bronstein et al. (2021), the composition of permutation invariant layers ($f=\psi\circ\phi\circ\psi\circ\phi\cdots$) allows the construction of functions $f$ that are equivariant to symmetry group actions. In the multigraph domain, this symmetry group includes permutations of both nodes and edges, as node permutations do not directly induce edge permutations due to the presence of parallel edges. The overall permutation equivariance of the MEGA-GNN model follows from the fact that each permutation invariant message passing layer operates independently on each node’s neighborhood, regardless of the ordering of nodes or edges. Specifically, for any permutation $g\in\sum_{n}$ acting on the set of node and edge indices, the model’s output satisfies $f(g\cdot X)=g\cdot f(X)$.
@@
Egressy et al. (2024) showed that a GNN can mimic a Breadth-First Search (BFS) algorithm to compute unique node IDs given pre-computed port numbers for the edges. We follow the same BFS-based approach and derive unique node ids without relying on pre-computed port numbers. Instead of the pre-computed port numbers, we use the unique edge labels provided by $L(e)$ to guide the node ID assignment process. As in Egressy et al. (2024), we use the Universal Approximation Theorem Hornik et al. (1989) for MLPs, to avoid explicit construction of the MEGA-GNN layers. We also assume that the MEGA-GNN aggregates the multi-edges by computing their minimum, which is followed by a node-level aggregation, where an MLP is applied element-wise to the incoming messages, followed by another minimum computation.
Following the approach of Egressy et al. (2024), the node ID assignment algorithm starts from a root node (also called the ego node) and assigns IDs to all the other nodes connected to it via message passing . We are not going to reiterate the setup of the entire proof and focus on the differences. Namely, instead of pre-computing port numbers, we use an edge labeling function, which assumes a strict total ordering among the edges based on the original edge features. In addition, the multi-edges are aggregated by selecting the edge with the minimum label.
Our MEGA-GNN model, which mimics Algorithm 1, assigns ids to each node connected to the root node. What remains to be shown is that those assigned node IDs are unique. First, note that nodes at different distances from the root cannot end up with the same node ID. A node at distance $k$ will receive its first proposal in round $k$ and, therefore, it will have an ID with exactly $k+1$ digits. Furthermore, an inductive argument shows that active nodes (nodes at the same distance) cannot have the same node IDs. Certainly, the argument is also true at the start when $X=\{r\}$. Now assuming all active nodes from the previous round $(k-1)$ had distinct node IDs, then the only way two active nodes (in round $k$) can have the same ID is if they accept a proposal from the same neighboring node. This is because, based on the induction hypothesis, proposals from different nodes will already differ in their first $k-1$ digits. If two active nodes accepted a proposal from the same node, then they would have received different edge labels —a strict total ordering among the edges enables assignment of distinct edge labels. In addition, because $m$ is added to all incoming labels, incoming labels cannot be the same as the outgoing labels. Therefore the active nodes always accept unique proposals.
@@ B.1 Hyperparameters
Table 5: Hyperparameter settings for AML and ETH datasets
@@ B.2 Multi-Edge Aggregation
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{MLP}\bigg{(}\text{SUM}\{\!\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\!\}\bigg{)}.\end{split}$ (19)
@@
In the PNA aggregation we combined [mean, min, max, std] statistics in combination with scalars [identity, amplification, attenuation] around the neighborhood of each artificial node, $N_{in}(v_{ij}^{art})=ME_{ij}$.
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{MLP}\bigg{(}\text{PNA}\{\!\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\!\}\bigg{)}.\end{split}$ (20)
@@
$\begin{split}\mathbf{h}_{ij}^{(l)}=\text{GenAgg}\bigg{(}\text{SUM}\{\!\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}\!\}\bigg{)}.\end{split}$ (21)
@@ Appendix C Additional Performance Metrics for AML Results
Table 6: Edge classification task results for AML Small datasets.
AML Small LI AML Small HI
@@
Table 7: Edge classification task results for AML Medium datasets.
AML Medium LI AML Medium HI
@@ Appendix E Ablation Experiment
In this section, we conduct additional experiments on the baseline method, Multi-GNN, to investigate the effect of random port numbering. The results, presented in Table 8, are obtained using the same set of hyperparameters as those used for MEGA-GIN and MEGA-PNA, which are applied to Multi-GIN and Multi-PNA, respectively.
Table 8: Multi-GNN ablations: The port numbering in Multi-GNN is based on a strict total ordering of the edges. In the permuted ablations, this port numbering is randomized to evaluate the impact of edge ordering on model performance.
Ablation AML Small HI AML Small LI AML Medium HI AML Medium LI ETH
Multi-GIN 62.65 $\pm$ 1.73 32.21 $\pm$ 0.99 67.72 $\pm$ 0.93 31.24 $\pm$ 2.12 51.34 $\pm$ 3.92
Multi-GIN (permuted) 61.36 $\pm$ 4.17 29.00 $\pm$ 3.49 67.58 $\pm$ 1.96 30.19 $\pm$ 1.49 49.13 $\pm$ 3.72
Multi-PNA 67.35 $\pm$ 2.89 35.39 $\pm$ 3.93 71.93 $\pm$ 1.54 43.81 $\pm$ 0.51 64.61 $\pm$ 1.40
Multi-PNA (permuted) 66.58 $\pm$ 1.86 33.05 $\pm$ 3.84 73.41 $\pm$ 1.38 42.61 $\pm$ 1.21 65.01 $\pm$ 0.97
Random assignment of port numbers leads to a slight performance drop for both models, with Multi-GIN experiencing a more significant decrease compared to Multi-PNA. Notably, Multi-PNA demonstrates greater robustness to permuted port numbers. Interestingly, on the ETH dataset, permuting the port numbers results in a small performance improvement for the PNA-based model, suggesting that the permutation equivariance property is less critical for this dataset.